OpenStudy (anonymous):

Find the volume of the solid generated by revolving the shaded region about the x-axis. 5x+2y =10

jigglypuff314 (jigglypuff314):

Hello and Welcome to OpenStudy! :) do you have a picture of where the shaded region would be?

OpenStudy (anonymous):

Hi, and thanks!

jigglypuff314 (jigglypuff314):

I'm guessing you're in calculus?

OpenStudy (anonymous):

Yes, Math 104

jigglypuff314 (jigglypuff314):

alright so if you rotate the shaded region around the x axis what do you think you would get? a washer? or cylinder thing?

OpenStudy (anonymous):

Couldn't you just calculate the area of the triangle, and multiply that by\[2\pi r\]to get the volume? I hate shaded regions

OpenStudy (anonymous):

Would that work Embryo? I hadn't thought of it... Hmm

OpenStudy (anonymous):

And to jigglypuff314, I have this so far:

OpenStudy (anonymous):

\[V \int\limits_{0}^{2} \pi (R(x))^2 dx\] \[R(x) = (10-5x)/2\] \[5x + 2(0) = 10\]

jigglypuff314 (jigglypuff314):

@Embryo I'm not getting the same answer as I would have if I did it the calculus way...

OpenStudy (anonymous):

Then dangit, i really hate shaded regions

jigglypuff314 (jigglypuff314):

jghost, the first two steps look right to me I'm not sure what the third thing you have is :/ just plug in what you have for R(x) into the integral and solve

OpenStudy (anonymous):

Sorry, the third thing was how I got the Upper Limit of Integration to be 2... but I'm having trouble solving :(

OpenStudy (anonymous):

@Embryo, revolving the region would give you a cone. You could use the formula for a cone's volume to avoid the integration method.

Parth (parthkohli):

Here's a non-calc way. By a little imagination, the volume of this thing is exactly the half of the volume of a cylinder with radius 5 and height 2.\[\pi \times 25 \times 2 \times \dfrac{1}{2}\]\[= 25\pi\]

Parth (parthkohli):

Is that the right answer...?

OpenStudy (anonymous):

\[V=\pi r^2\frac{ h }{ 3 }\]where r is the radius, being 5, and h= the height from the base to the tip, which is 2, like that @SithsAndGiggles ?

jigglypuff314 (jigglypuff314):

@Embryo yes :) that is the answer I got from the integration way

OpenStudy (anonymous):

@Embryo, yeah

OpenStudy (anonymous):

So @JGhost you now have multiple ways of solving this, pick your favorite and roll with it

jigglypuff314 (jigglypuff314):

\[\large \pi \int\limits_{0}^{2} (-2.5x + 5)^2~~ dx\] would be the integration way, which you can/should just plug into calculator :P

OpenStudy (anonymous):

Hmm I went with the calc method and got 50pi/3 as the ans?

Parth (parthkohli):

OK, then go with the calc method.

jigglypuff314 (jigglypuff314):

correct! :)

Parth (parthkohli):

This makes way for another kind of curiosity. Why was the non-calc way wrong?

Parth (parthkohli):

Ah, I see... never mind.

OpenStudy (anonymous):

Sweet! The system accepted it as correct! Thanks peeps. @jigglypuff314 @Embryo @ParthKohli @SithsAndGiggles

jigglypuff314 (jigglypuff314):

glad we could help ^_^